# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

### v3

@@ 3.3. Setting up the bootstrap argument
which respectively control curvature energies, renormalized entropy for the transverse-traceless ($TT$) second fundamental form, and point-wise norm of gauge variable and $TT$ tensor $\Sigma$. These norms are defined in the rescaled setting associated with the dynamical time variable $T\in[T_{0},T_{\infty})$.
@@ 3.4. Estimates for the metric components
where $N$ is the lapse function, $\Sigma_{ij}$ is the transverse-traceless part of the second fundamental form, $\tau$ is the mean curvature, and $\varphi$ is a smooth weight function. For a fixed tangent vector $Y\in T_{x}M$, we obtain from (171)
@@
Propositions 3.1 and 3.2 establish uniform two-sided bounds on the evolving spatial metric $g(T,x)$ in terms of the initial metric $g_{0}:=g(T_{0},x)$, valid on the entire bootstrap interval $T\in[T_{0},T_{\infty}]$, with $T_{0}\gg 1$ sufficiently large. More precisely, there exists a numerical constant $C\geq 1$, such that
@@ 3.5. Controlling the Sobolev Constants
form a foliation by compact $2$-dimensional Riemannian manifolds (with induced metric still denoted $g(T)$). Assume the initial data on $M_{T_{0}}$ satisfy the hypotheses of Section 2 and that the bootstrap assumption (2.10) holds along the entire slab $T\in[T_{0},T_{\infty})$. Then for every $T\in[T_{0},T_{\infty})$ the isoperimetric constant satisfies
@@
where $\alpha(T),\beta(T)$ are likewise controlled by 3.2. Integrating (107) over $U_{0}$ and $U^{c}_{0}$ yields
@@
$\|\psi\|_{L^{\infty}}=\|f\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big({\rm Vol}_{g}(M)^{\frac{1}{3}}\|\nabla f\|_{L^{4}}+{\rm Vol}_{g}(M)^{-\frac{1}{4}}\|f\|_{L^{4}}\Big).$
@@
$\|\psi\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big({\rm Vol}_{g}(M)^{\frac{1}{3}}\|\nabla\psi\|_{L^{4}}+{\rm Vol}_{g}(M)^{-\frac{1}{4}}\|\psi\|_{L^{4}}\Big),$
@@
For the dynamical manifold $(M_{T},g(T))$, the Sobolev inequalities 109-110 hold with the constant $C\bigg(\max(1,I(M,g))\bigg)$ being uniformly bounded by $O(1)$ purely numerical entity $\forall T\in[T_{0},T_{\infty})$ in light of the proposition $3.$
Lastly, we note that on $(M,g)$ we can obtain compactness theorems for Sobolev embeddings. This can be proven by working on local charts using the uniform estimates on the metric $g(t)$ on the interval $T_{0},T_{\infty}]$ in light of propositions 3.1-3.2 and gluing the charts together in a compatible manner. In particular, on $(M,g)$ for a $C^{\infty}$ tensor field $\psi$, the Sobolev norm $H^{I}(M)$ is defined as
@@ 4. Estimates
(115) $\operatorname{Ric}_{ij}=\mathfrak{T}_{ij}+\frac{1}{n}R(g)g_{ij},$
@@ 4.1. Elliptic estimates
(118) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
@@
Let $(M,g)$ be a compact Riemannian $3$-manifold with metric $g$ satisfying the bootstrap assumption (87), and let $\Sigma$ denote a symmetric transverse-traceless (TT) $2$-tensor on $M$. Consider the lapse function $N$ solving the elliptic equation
$\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|_{g}^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|_{g}^{2},$
@@
(128) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
Let us denote the operator $\Delta_{g}+(|\Sigma|^{2}+\frac{1}{n})$ by $\mathscr{L}$. We claim that an estimate of the following type holds for $u\in C^{\infty}(M)$
@@
$\mathscr{L}\nabla^{I}(\frac{N}{n}-1)=-\Delta_{g}\nabla^{I}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})\nabla^{I}(\frac{N}{n}-1)=\nabla^{I}\mathscr{L}(\frac{N}{n}-1)$
@@
(133) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
@@
(134) $\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)^{2}\mu_{g}=-\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)\mu_{g}.$
@@
Therefore, 134 reads
$\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)^{2}\mu_{g}+(\frac{1}{n}-\frac{1}{n+1})\int_{M}(\frac{N}{n}-1)^{2}\mu_{g}\leq(n+1)\int_{M}|\Sigma|^{4}\mu_{g}$
which by means of the inequality 3.4 yields
@@
Now iterate this using 132 to obtain
@@
4.2. Estimating the lower order norm of TT tensor $\Sigma$ via transport equation for $\Sigma$
In this section, we prove that the weighted lower order norm of the $TT$ tensor $\Sigma$ is controlled by the top order norm of $(\mathfrak{T},\Sigma)$ and the initial data. Recall the equation for $\Sigma$
@@ 4.2. Estimating the lower order norm of TT tensor $\Sigma$ via transport equation for $\Sigma$
Theorem 4.1 (Coercive control of the reversed entropy).
@@
and $\mathcal{O}$ be the top order norm 84 Then the following estimate for the re-scaled reversed entropy holds
@@
$\mathcal{F}\lesssim 1+\mathcal{I}^{0}+\mathcal{O}$ and $\mathcal{N}^{\infty}\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2}$ by elliptic estimates 4.1.
@@ 4.3. Energy Estimates
Theorem 4.2 (Top-order energy bound).
@@
The maximum derivative on lapse function in the previous estimate is $I+1$ and therefore by elliptic estimate, it is bounded by $\Sigma||^{2}_{H^{I-1}(M)}$ which yields $e^{-2\gamma t}$ decay. The next term is estimated as
@@
Here, the elliptic estimate 4.1 for the lapse function is used. The next term is estimated as
@@
Now note that the first term in the expression involving $P^{I}$ contains term $I$ which is negative definite. Therefore, the collection of every term yields
@@ 5. Proof of the main results
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on MM of pointwise constant negative scalar curvature, in the sense that
@@
Let $N$ denote the rescaled lapse in the CMC gauge so that the spacetime metric reads
@@
We emphasize that in general, the trace-free renormalized Ricci tensor $\mathfrak{T}$ does not decay to zero as $T\to\infty$. This indicates that the limiting metric $\widetilde{g}:=\lim_{T\to\infty}g(T)$ fails to be Einstein, even though it has constant scalar curvature. The underlying obstruction is geometric: the compact manifold MM may not admit any Einstein metric (e.g., hyperbolic metric in the case $n=3$), and hence the evolution governed by the Einstein equations with $\Lambda>0$ does not necessarily drive the geometry toward an Einstein configuration.
@@
By the first variation formula for volume under the flow, and using the uniform exponential decay estimates for the lapse function $N$ (cf. Section 4.3), we obtain
@@
with lapse $N$ uniformly bounded above by the estimates $\|\frac{N}{n}-1\|_{L^{\infty}(M)}\lesssim e^{-2T},~T\gg 1$. Causality implies
@@
In light of the estimates of theorem 1.1 and the scaling property of 59 in section 3, the physical metric $\widetilde{g}(T)\approx\varphi^{-2}g(T)$ for the re-scaled nonphysical metric $g(T)$. Hence,
@@
Note that for the pure vacuum i.e., $\Lambda=0$, one has $\varphi=\tau$ and therefore the right hand side does not have a decay structure. For $\Lambda>0$, one observes $|\varphi(T)|\lesssim e^{-T}$ and $|\tau|\lesssim\sqrt{e^{-2T}+3\Lambda}=\sqrt{e^{-2T}+3n(n+1)}\gtrsim\sqrt{n(n+1)}$ since we fixed $\Lambda=n(n+1)$ in our framework. The length of the spatial projection $\pi_{M}\circ\gamma$ of $\gamma$ in $(M,g(T))$ satisfies

### v4

@@ 5.1. Setting up the bootstrap argument
which respectively control curvature energies, renormalized entropy for the transverse-traceless (TTTT) second fundamental form, and point-wise norm of gauge variable and TTTT tensor $\Sigma$. These norms are defined in the rescaled setting associated with the dynamical time variable $T\in[T_{0},T_{\infty})$.
@@ 5.2. Estimates for the metric components
where NN is the lapse function, $\Sigma_{ij}$ is the transverse-traceless part of the second fundamental form, $\tau$ is the mean curvature, and $\varphi$ is a smooth weight function. For a fixed tangent vector $Y\in T_{x}M$, we obtain from (250)
@@
Propositions 5.1 and 5.2 establish uniform two-sided bounds on the evolving spatial metric $g(T,x)$ in terms of the initial metric $g_{0}:=g(T_{0},x)$, valid on the entire bootstrap interval $T\in[T_{0},T_{\infty}]$, with $T_{0}\gg 1$ sufficiently large. More precisely, there exists a numerical constant $C\geq 1$, such that
@@ 5.3. Controlling the Sobolev Constants
form a foliation by compact 22-dimensional Riemannian manifolds (with induced metric still denoted $g(T)$). Assume the initial data on $M_{T_{0}}$ satisfy the hypotheses of Section 2 and that the bootstrap assumption (2.10) holds along the entire slab $T\in[T_{0},T_{\infty})$. Then for every $T\in[T_{0},T_{\infty})$ the isoperimetric constant satisfies
@@
where $\alpha(T),\beta(T)$ are likewise controlled by 5.2. Integrating (186) over $U_{0}$ and $U^{c}_{0}$ yields
@@
$\|\psi\|_{L^{\infty}}=\|f\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big(Vol_{g}(M)^{\frac{1}{3}}\|\nabla f\|_{L^{4}}+{Vol}_{g}(M)^{-\frac{1}{4}}\|f\|_{L^{4}}\Big).$
@@
$\|\psi\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big({Vol}_{g}(M)^{\frac{1}{3}}\|\nabla\psi\|_{L^{4}}+{Vol}_{g}(M)^{-\frac{1}{4}}\|\psi\|_{L^{4}}\Big),$
@@
For the dynamical manifold $(M_{T},g(T))$, the Sobolev inequalities 188-189 hold with the constant $C\bigg(\max(1,I(M,g))\bigg)$ being uniformly bounded by $O(1)$ purely numerical entity $\forall T\in[T_{0},T_{\infty})$ in light of the proposition 5.3.
Lastly, we note that on $(M,g)$ we can obtain compactness theorems for Sobolev embeddings. This can be proven by working on local charts using the uniform estimates on the metric $g(t)$ on the interval $T_{0},T_{\infty}]$ in light of propositions 5.1-5.2 and gluing the charts together in a compatible manner. In particular, on $(M,g)$ for a $C^{\infty}$ tensor field $\psi$, the Sobolev norm $H^{I}(M)$ is defined as
@@
(194) $\operatorname{Ric}_{ij}=\mathfrak{T}_{ij}+\frac{1}{3}R(g)g_{ij},$
@@ 5.4. Elliptic estimates
(197) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{3})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
@@
Let $(M,g)$ be a compact Riemannian 33-manifold with metric $g$ satisfying the bootstrap assumption (166), and let $\Sigma$ denote a symmetric transverse-traceless ($T$) 22-tensor on $M$. Consider the lapse function NN solving the elliptic equation
$\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|_{g}^{2}+\frac{1}{3}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|_{g}^{2},$
@@
(207) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{3})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
Let us denote the operator $\Delta_{g}+(|\Sigma|^{2}+\frac{1}{3})$ by $\mathscr{L}$. We claim that an estimate of the following type holds for $u\in C^{\infty}(M)$
@@
$\mathscr{L}\nabla^{I}(\frac{N}{n}-1)=-\Delta_{g}\nabla^{I}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{3})\nabla^{I}(\frac{N}{n}-1)=\nabla^{I}\mathscr{L}(\frac{N}{n}-1)$
@@
(212) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{3})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
@@
(213) $\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}(|\Sigma|^{2}+\frac{1}{3})(\frac{N}{n}-1)^{2}\mu_{g}=-\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)\mu_{g}.$
@@
Therefore, 213 reads
$\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)^{2}\mu_{g}+(\frac{1}{3}-\frac{1}{n+1})\int_{M}(\frac{N}{n}-1)^{2}\mu_{g}\leq(n+1)\int_{M}|\Sigma|^{4}\mu_{g}$
which by means of the inequality 5.4 yields
@@
Now iterate this using 211 to obtain
@@
$5.$ Estimating the lower order norm of $T$ tensor $\Sigma$ via transport equation for $\Sigma$
In this section, we prove that the weighted lower order norm of the TTTT tensor $\Sigma$ is controlled by the top order norm of $(\mathfrak{T},\Sigma)$ and the initial data. Recall the equation for $\Sigma$
@@
Proposition 5.7 (Coercive control of the reversed entropy).
@@
and $\mathcal{O}$ be the top order norm 91 Then the following estimate for the re-scaled reversed entropy holds
@@
$\mathcal{F}\lesssim 1+\mathcal{I}^{0}+\mathcal{O}$ and $\mathcal{N}^{\infty}\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2}$ by elliptic estimates 5.6.
@@ 5.6. Energy Estimates
Proposition 5.9 (Top-order energy bound).
@@
The maximum derivative on the lapse function in the previous estimate is $I+1$ and therefore by the elliptic estimate, it is bounded by $\Sigma||^{2}_{H^{I-1}(M)}$ which yields $e^{-2\gamma t}$ decay. The next term is estimated as
@@
Here, the elliptic estimate 5.6 for the lapse function is used. The next term is estimated as
@@
Now note that the first term in the expression involving $P^{I}$ contains term II which is negative definite. Therefore, the collection of every term yields
@@ 6. Proof of the main results
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on $M$ of pointwise constant negative scalar curvature, in the sense that
@@
Let NN denote the rescaled lapse in the CMC gauge so that the spacetime metric reads
@@
We emphasize that in general, the trace-free renormalized Ricci tensor $\mathfrak{T}$ does not decay to zero as $T\to\infty$. This indicates that the limiting metric $\widetilde{g}:=\lim_{T\to\infty}g(T)$ fails to be Einstein, even though it has constant scalar curvature. The underlying obstruction is geometric: the compact manifold $M$ may not admit any Einstein metric (e.g., hyperbolic metric in the case $n=3$), and hence the evolution governed by the Einstein equations with $\Lambda>0$ does not necessarily drive the geometry toward an Einstein configuration.
@@
By the first variation formula for volume under the flow, and using the uniform exponential decay estimates for the lapse function NN (cf. Section 5.6), we obtain
@@
with lapse NN uniformly bounded above by the estimates $\|\frac{N}{n}-1\|_{L^{\infty}(M)}\lesssim e^{-2T},~T\gg 1$. Causality implies
@@
In light of the estimates of theorem 1.1 and the scaling property of 66 in section 3, the physical metric $\widetilde{g}(T)\approx\varphi^{-2}g(T)$ for the re-scaled nonphysical metric $g(T)$. Hence,
@@
Note that for the pure vacuum i.e., $\Lambda=0$, one has $\varphi=\tau$ and therefore the right hand side does not have a decay structure. For $\Lambda>0$, one observes $|\varphi(T)|\lesssim e^{-T}$ and $|\tau|=\sqrt{e^{-2T}+3\Lambda}=\sqrt{e^{-2T}+3n(n+1)}$ since we fixed $\Lambda=n(n+1)$ in our framework. The length of the spatial projection $\pi_{M}\circ\gamma$ of $\gamma$ in $(M,g(T))$ satisfies